OpenStudy (anonymous):

Please help! I don't get this at all... If and ∠ 1 ≅ ∠ 2, prove that ∠ NLO ≅ ∠ NPM. Where is Hector's mistake and how must he correct it?

OpenStudy (anonymous):

Statement 1 because the information in the proof is not given in the problem. Statement 3 because the angle addressed is irrelevant to the proof. Statement 4 because the triangles are inaccurately labeled. Statement 5 because those angles are not corresponding parts.

OpenStudy (anonymous):

Statements Reasons 1. LN ~ NP Given 2. ∠ 1 ≅ ∠ 2 Given 3. ∠ Q ≅ ∠ Q Reflexive Property 4. ΔLNO ≅ ΔPNM Angle-Angle-Side Postulate 5. ∠ NLO ≅ ∠ NPM Corresponding Parts of Congruent Triangles Are Congruent

OpenStudy (anonymous):

Well, we can say: \[ \angle 1 = \angle 2 = \alpha \] Now for both triangle △NLO we can say: \[ \angle L = 180^\circ - ( \angle N + \angle 2) = 180^\circ - (\angle N + \alpha) \] and for triangle and △NPM we can say: \[ \angle P = 180^\circ - ( \angle N + \angle 1) = 180^\circ - (\angle N + \alpha) \] Therefore ∠L≅∠P And for the other part, first you didn't say anything about LN or NP so I would say A for using information that is not given in the problem. He also uses angle ∠Q that is in neither one of those triangles in his triangle congruence claim. So I would say that B for addressing irrelevant angles to the proof.